# Extremal Steklov-Neumann eigenvalues

### v1

@@ Acknowledgments
    $u(x)=\sum_{k\in\mathbb{N}}c_{k}\sigma_{k}^{-\frac{1}{2}}U_{k}(x)$ (12)
@@
    $\lambda_{1}^{-1}(\rho)=\max_{c\in\ell^{2}(\mathbb{N})}\frac{c^{t}A(\rho)c}{c^{t}c}.$ (13)
    Furthermore, the map $\rho\mapsto\lambda_{k}^{-1}(\rho)$ is convex for $k=1$.
@@
$T_{ij}(\rho)=\frac{1}{\sqrt{\sigma_{i}\sigma_{j}}}\int_{\Gamma}\rho\phi_{i}\phi_{j}\,ds=A_{ij}(\rho).$ (14)
@@
(b) We follow the proof of [10, Prop. 2.2]. Suppose $\rho_{n}\overset{\ast}{\rightharpoonup}\rho$ in $L^{\infty}(\Gamma)$ and $\lambda_{k}^{(n)}$ is the $k$-th eigenvalue satisfying (4) for density $\rho_{n}$ with an associated eigenfunction $u_{k}^{(n)}$, normalized so that
$\int_{\Gamma}\rho_{n}\left(u_{k}^{(n)}\right)^{2}ds=1\qquad\textrm{ and }\qquad\lambda_{k}^{(n)}=\int_{\Omega}\left\lvert\nabla u_{k}^{(n)}\right\rvert^{2}dx.$ (15)
@@
(c) For fixed $c\in\ell^{2}(\mathbb{N})$, we see that the map $\rho\mapsto c^{t}A(\rho)c$ is linear. Using the variational principle (13), $\rho\mapsto\lambda_{1}^{-1}(\rho)$ is the pointwise supremum of linear functions of $\rho$ and hence convex.
@@
$\lambda_{\varepsilon}$ $=\lambda+\varepsilon\nu+\mathcal{O}(\varepsilon^{2}),$ (16a)
$u_{\varepsilon}$ $=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\Big(\delta_{k,k_{0}}a_{j}+\varepsilon b_{j,k}+\mathcal{O}(\varepsilon^{2})\Big)u_{j,k},$ (16b)
@@
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}b_{j,k}\rho(\lambda_{k}-\lambda_{k_{0}})u_{j,k}$ $=\sum_{j=1}^{m_{k_{0}}}\left(\lambda_{k_{0}}\delta\rho+\nu\rho\right)a_{j}u_{j,k_{0}}.$ (17)
@@
Using the direct method of the calculus of variations, the existence of a minimizer in (7) and a maximizer in (8) follows from (9), Lemma 2.1(c), and Lemma 2.2(b).
For $k=1$, the optimization problem in (7) is equivalent to $\max_{\rho\in\mathcal{A}_{\alpha}}\ \lambda^{-1}_{1}(\rho)$. A bang-bang minimizer exists following from Lemma 2.1(a), Lemma 2.2(c), and the fact that a bounded convex function on a compact convex set attains its maximum at an extreme point. ∎
@@
Since $\rho^{\bigtriangledown}$ is a local minimizer, $M(\delta\rho)$ must be positive definite for every admissible $\delta\rho$, i.e.,
$\int_{\Gamma}q_{a}(s)\delta\rho(s)\,ds\leq 0,\qquad\quad\forall a\in\mathbb{R}^{m_{k}},\,|a|=1,$ (18)
@@
$A_{0}\coloneqq\left\{\rho_{k}^{\bigtriangledown}=0\right\},\qquad A_{1}\coloneqq\left\{\rho_{k}^{\bigtriangledown}=1\right\},\quad\textrm{and}\quad A\coloneqq\left\{0<\rho_{k}^{\bigtriangledown}<1\right\}.$
@@
Proof of Theorem 1.5..
We prove the statement for rotation symmetry; the statement for reflection symmetry can similarly be proven. Suppose $\Omega$ has a $\frac{2\pi}{N}$-rotational symmetry, $R_{N}\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$, so that $R_{N}\Omega=\Omega$. Let $\rho_{1}^{\triangle}$ be a maximizer and consider the admissible density, $\overline{\rho}(s)=\frac{1}{N}\sum_{j\in[N]}\rho_{1}^{\triangle}(R^{N}s)$. By the convexity of $\rho\mapsto\lambda_{1}^{-1}(\rho)$ (see Lemma 2.2(c)), Jensen’s inequality, and rotational symmetry of $\lambda_{1}$, we then have that
$\lambda_{1}^{-1}(\overline{\rho})\leq\frac{1}{N}\sum_{j\in[N]}\lambda_{1}^{-1}\left(\rho_{1}^{\triangle}\circ R^{N}\right)=\lambda_{1}^{-1}(\rho_{1}^{\triangle}).$
This shows that $\lambda_{1}(\overline{\rho})\geq\lambda_{1}(\rho_{1}^{\triangle})$ and since $\rho_{1}^{\triangle}$ is a maximizer, the rotational-symmetric density $\overline{\rho}$ is also a maximizer. ∎
@@ 3. Analysis of extremal eigenvalues for a disk
3.1. Proof of Theorem 1.8(1)
Proof.
Let $\rho$ be an admissible density in $\mathcal{A}_{\alpha}$ and write its average value $\overline{\rho}$ as the convex combination of rotations of $\rho$, $\overline{\rho}(\theta)=\frac{1}{2\pi}\int_{0}^{2\pi}\rho(\theta-\theta_{0})\,d\theta_{0}=\frac{1}{2\pi}\int_{0}^{2\pi}\rho(\theta_{0})\,d\theta_{0}=\alpha$. By the convexity of $\rho\mapsto\lambda_{1}^{-1}(\rho)$ (see Lemma 2.2(c)), Jensen’s inequality, and rotational symmetry of $\lambda_{1}$, we then have that
$\lambda_{1}^{-1}(\overline{\rho})\leq\frac{1}{2\pi}\int_{0}^{2\pi}\lambda_{1}^{-1}\left(\rho(\cdot-\theta_{0})\right)d\theta_{0}=\lambda_{1}^{-1}(\rho).$
This shows that $\lambda_{1}(\overline{\rho})\geq\lambda_{1}(\rho)$ and since $\rho$ is an arbitrary admissible density, the constant density $\overline{\rho}=\alpha$ is a maximizer for (8) for $k=1$. ∎
Remark 3.1.
Note that if $\rho\mapsto\lambda_{k}^{-1}(\rho)$ were convex for $k\geq 2$, then the proof of Theorem 1.8(1) would also apply to $\lambda_{k}^{-1}$. Since, by Theorem 1.8(2), the constant density is not the maximizer for even $k\geq 2$, this would be a contradiction; we conclude that $\rho\mapsto\lambda_{k}^{-1}(\rho)$ is not convex for $k\geq 2$.
@@
$M=mI+\frac{1}{2\pi\varepsilon}\sum_{j\in J}\sin(m\left\lvert S_{j}\right\rvert)\begin{pmatrix}\cos(2m\theta_{j})&\sin(2m\theta_{j})\\ \sin(2m\theta_{j})&-\cos(2m\theta_{j})\end{pmatrix}$. (19)
@@ 4.1. Computation of weighted Steklov eigenvalues
To solve the weighted Steklov eigenproblem (4) computationally, we first reformulate it using layer potential methods and then discretize the resulting integral equation to obtain a matrix eigenvalue problem. The eigenfunction $u(x)$ is represented using a modified single layer potential, $\varphi$,
$u(x)=\int_{\Gamma}\Phi(x-y)(\varphi(y)-\overline{\varphi})\,ds(y)+\overline{\varphi},$ (20)
where $\Phi(x)=\frac{1}{2\pi}\log\left\lvert x\right\rvert$ and $\overline{\varphi}=\int_{\Gamma}\varphi(y)\,ds(y)/\int_{\Gamma}\,ds(y)$. The modification ensures uniqueness of the solution. Taking into account well-known expressions (see e.g., [20]) for the jump of the single layer potential and its normal derivative across $\Gamma$, the weighted Steklov eigenproblem (4) reduces to the integral eigenvalue equation for $(\lambda,\varphi)$,
$A[\varphi]=\lambda\rho B[\varphi].$ (21)
@@
where $\tt A$, $\tt B$, and $\tt X$ are discretizations of $A$, $B$, and $\varphi$, respectively. The diagonal matrix $\tt R$ is the discretizations of $\rho$. More details about this method can be found in [1]. We implement these methods in MATLAB using MPSpack [7].
@@ 4.2. Computation of extremal Steklov eigenvalues
This minimax problem can be numerically solved using a sequential quadratic programming (SQP) method [9], as implemented in MATLAB’s fminimax function. The gradient of the (numerically simple) eigenvalues is provided using Lemma 2.2(d). We use as convergence criteria a tolerance on either the relative change on stepsize or function value. The maximization problem (8) is analogously reformulated as a maximin problem and solved using the same methods.
@@ 5.1. Minimizers for the disk
Interestingly, $\rho_{k}^{\bigtriangledown}$ is a bang-bang function for all $k\geq 1$, even though we were only able to prove this for $k=1$ (see Theorem 1.1).
@@
Figure 4. (Left) Maximizing densities, $\rho_{k}^{\triangle}$, and associated eigenfunctions for (8) with $\Omega=\mathbb{D}$, $\alpha=0.5$, and even $k=2,4,6,8$ (values indicated). The red part of the boundary is where $\rho=1$, i.e., where the Steklov boundary condition is imposed. Interestingly, the eigenfunctions are constant on intervals where $\rho\neq 1$. (Right) The eigenfunctions are plotted on $\Omega=\mathbb{D}$. See Section 5.2.
@@
Figure 5. Solutions for (8) with $\Omega=\mathbb{D}$, $\alpha=0.5$, and $k=1,2,3,4$ (values indicated). In the left hand panels, the maximizing density $\rho_{k}^{\triangle}$ is plotted in blue and the corresponding eigenfunction(s) are plotted in red. For $k$ even, the eigenvalue is simple and the right hand panel shows that the optimality condition (11) is satisfied. For $k$ odd, the eigenvalue has multiplicity 22 and the right hand panels show that the optimality condition (11) is satisfied for the two eigenfunctions plotted in the left hand panel. See Section 5.2.
5.2. Maximizers for the disk
For $k\geq 2$, the maximizer $\rho_{k}^{\triangle}$ in (8) has a non-trivial form shown in Figures 4 and 5. It is the $k$-repetition of an interval that is constant (=1=1) on a first subinterval and is strictly positive and convex on a second subinterval.
In Figure 6, we plot $\alpha$ vs. $\alpha\lambda_{2,\alpha}^{\triangle}$. This indicates that, for $k$ even, except in the limit that $\alpha\to 0$, we have that $\lambda_{2,\alpha}^{\triangle}<\frac{2}{\alpha}$, implying that $\lambda_{k,\alpha}^{\triangle}=\frac{k}{2}\lambda_{2,\alpha}^{\triangle}<\frac{k}{\alpha}=\frac{2\pi}{\alpha\left\lvert\Gamma\right\rvert}k$, which shows that the Hersch-Payne-Schiffer inequality [16, 12] is not strict for the extremal weighted Steklov eigenvalue problem (and hence the extremal S-N eigenvalue problem) for $k$ even and $\alpha>0$.
We remark that the oscillations for the maximizing densities, $\rho_{k}^{\triangle}$, are reminiscent of the maximizing domains with oscillatory boundaries for the shape optimization problem studied in [1].
@@ 5.2. Maximizers for the disk
$\frac{\lambda_{k,\alpha}^{\bigtriangledown}}{\lambda_{k,\alpha}^{W}}\to\alpha\sigma_{1,\alpha}^{\bigtriangledown}=\alpha\sigma_{1}(\mathcal{I}_{\alpha}^{2})\qquad\qquad\textrm{and}\qquad\qquad\frac{\lambda_{k,\alpha}^{\triangle}}{\lambda_{k,\alpha}^{W}}\to\alpha\lambda_{2,\alpha}^{\triangle}.$ (22)
In Figure 6, we plot $\alpha$ vs. $\alpha\lambda_{1,\alpha}^{\bigtriangledown}=\alpha\sigma_{1,\alpha}^{\bigtriangledown}$ and $\alpha\lambda_{2,\alpha}^{\triangle}$. We see that for $\alpha$ small, say $\alpha=0.02$, we have that $\alpha\sigma_{1,\alpha}^{\bigtriangledown}\approx 0.2$ which indicates that $\sigma_{k}(\mathcal{I}_{\alpha}^{k+1})$ is only $20\%$ what Weyl’s law would predict! Furthermore, for $\alpha$ small, we have that $\alpha\lambda_{2,\alpha}^{\triangle}\approx 2$ which indicates that $\lambda_{k,\alpha}^{\triangle}$ is twice what Weyl’s law would predict for large $k$. For other examples of where extremal eigenvalues are a multiplicative factor different than predicted by Weyl’s law, see [12, 1, 17, 19, 18].
@@ Appendix A A counterexample to convexity properties of higher eigenvalues
with $0\leq\alpha<\beta$. In [15, Theorem 9.1.3], it is claimed that $\rho\mapsto\lambda^{-1}_{k}(\rho)$ is convex for fixed $k\geq 1$, but this is not true for $k\geq 2$ (as the author agreed in personal communication).
@@
$2\tan(\sqrt{\lambda})=-\tan\left(2\sqrt{\lambda}\right)=\frac{-2\tan(\sqrt{\lambda})}{1-\tan^{2}(\sqrt{\lambda})},$ (23)
@@
$f_{t}\big|_{t=0,\,\lambda=\eta_{k}}=-\frac{3}{4}\left(-2\tan(2\sqrt{\eta_{k}})+\sqrt{\eta_{k}}\sec^{2}(2\sqrt{\eta_{k}})+\tan(\sqrt{\eta_{k}})-4\sqrt{\eta_{k}}\sec^{2}(\sqrt{\eta_{k}})\right).$ (24)
@@
Finally, we remark that convexity/concavity properties for higher eigenvalues of other operators is also misstated, e.g., [15, Theorem 8.1.2] for the Schrödinger equation and [15, Theorem 10.1.1] for the conductivity equation.

### v2

@@ Acknowledgments
    (12) $u(x)=\sum_{k\in\mathbb{N}}c_{k}\sigma_{k}^{-\frac{1}{2}}U_{k}(x)$
@@
    (13) $\lambda_{1}^{-1}(\rho)=\max_{c\in\ell^{2}(\mathbb{N}),\,(c,b)=0}\frac{c^{t}A(\rho)c}{c^{t}c},$
    where $b=\{b_{k}\}_{k\in\mathbb{N}}\in\ell^{2}(\mathbb{N})$ is given by $b_{k}=\int_{\Gamma}\sigma_{k}^{-1/2}\phi_{k}(s)\rho(s)\,ds$.
@@
(14) $T_{ij}(\rho)=\frac{1}{\sqrt{\sigma_{i}\sigma_{j}}}\int_{\Gamma}\rho\phi_{i}\phi_{j}\,ds=A_{ij}(\rho).$
@@
(b) We follow the proof of [undefi, Prop. 2.2]. Suppose $\rho_{n}\overset{\ast}{\rightharpoonup}\rho$ in $L^{\infty}(\Gamma)$ and $\lambda_{k}^{(n)}$ is the $k$-th eigenvalue satisfying (4) for density $\rho_{n}$ with an associated eigenfunction $u_{k}^{(n)}$, normalized so that
(15) $\int_{\Gamma}\rho_{n}\left(u_{k}^{(n)}\right)^{2}ds=1\qquad\textrm{ and }\qquad\lambda_{k}^{(n)}=\int_{\Omega}\left\lvert\nabla u_{k}^{(n)}\right\rvert^{2}dx.$
@@
(c) The variational characterization (13) follows from part (a).
@@
(16a) $\lambda_{\varepsilon}$ $=\lambda+\varepsilon\nu+\mathcal{O}(\varepsilon^{2}),$
(16b) $u_{\varepsilon}$ $=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\Big(\delta_{k,k_{0}}a_{j}+\varepsilon b_{j,k}+\mathcal{O}(\varepsilon^{2})\Big)u_{j,k},$
@@
(17) $\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}b_{j,k}\rho(\lambda_{k}-\lambda_{k_{0}})u_{j,k}$ $=\sum_{j=1}^{m_{k_{0}}}\left(\lambda_{k_{0}}\delta\rho+\nu\rho\right)a_{j}u_{j,k_{0}}.$
@@
Using the direct method of the calculus of variations, the existence of a minimizer in (7) and a maximizer in (8) follows from (9), Lemma 2.1(c), and Lemma 2.2(b). ∎
@@
Since $\rho_{k}^{\bigtriangledown}$ is a local minimizer, $M(\delta\rho)$ must be positive definite for every admissible $\delta\rho$, i.e.,
(18) $\int_{\Gamma}q_{a}(s)\delta\rho(s)\,ds\leq 0,\qquad\quad\forall a\in\mathbb{R}^{m_{k}},\,|a|=1,$
@@
$A_{0}\coloneqq\left\{\rho_{k}^{\bigtriangledown}=0\right\},\quad A_{1}\coloneqq\left\{\rho_{k}^{\bigtriangledown}=1\right\},\quad\textrm{and}\quad A\coloneqq\left\{0<\rho_{k}^{\bigtriangledown}<1\right\}.$
@@ 3. Analysis of extremal eigenvalues for a disk
(19) $M=mI+\frac{1}{2\pi\varepsilon}\sum_{j\in J}\sin(m\left\lvert S_{j}\right\rvert)\begin{pmatrix}\cos(2m\theta_{j})&\sin(2m\theta_{j})\\ \sin(2m\theta_{j})&-\cos(2m\theta_{j})\end{pmatrix}$.
@@ 4.1. Computation of weighted Steklov eigenvalues
To solve the weighted Steklov eigenproblem (4) computationally, we first reformulate it using layer potential methods and then discretize the resulting integral equation to obtain a matrix eigenvalue problem. In this section, we additionally assume that the domain $\Omega$ is simply connected with $C^{2}$ boundary. The eigenfunction $u(x)$ is represented using a modified single layer potential, $\varphi$,
(20) $u(x)=\int_{\Gamma}\Phi(x-y)(\varphi(y)-\overline{\varphi})\,ds(y)+\overline{\varphi},$
where $\Phi(x)=\frac{1}{2\pi}\log\left\lvert x\right\rvert$ and $\overline{\varphi}=\int_{\Gamma}\varphi(y)\,ds(y)/\int_{\Gamma}\,ds(y)$. The modification ensures uniqueness of the solution. Taking into account well-known expressions (see e.g., [undefv]) for the jump of the single layer potential and its normal derivative across $\Gamma$, the weighted Steklov eigenproblem (4) reduces to the integral eigenvalue equation for $(\lambda,\varphi)$,
(21) $A[\varphi]=\lambda\rho B[\varphi].$
@@
where $\tt A$, $\tt B$, and $\tt X$ are discretizations of $A$, $B$, and $\varphi$, respectively. The diagonal matrix $\tt R$ is the discretizations of $\rho$. More details about this method can be found in [undef]. We implement these methods in MATLAB using MPSpack [undeff]. An accuracy test in [undef] demonstrated that an error less than $10^{-6}$ could be achieved with 512 points for a constant density.
@@ 4.2. Computation of extremal Steklov eigenvalues
This minimax problem can be numerically solved using a sequential quadratic programming (SQP) method [undefh], as implemented in MATLAB’s fminimax function. The gradient of the (numerically simple) eigenvalues is provided using Lemma 2.2(d). We use as convergence criteria a tolerance on either the relative change on stepsize or function value. The maximization problem (8) is analogously reformulated as a maximin problem and solved using the same methods.
In this paper, we focus on the numerical implementation of the first few eigenvalues, e.g., $k\leq 12$. For moderately large values of $k$, the discretization must be refined to resolve high frequency rapid oscillations. In addition, techniques such as tracking only a subset of eigenvalues and using continuation of subspace methods may be more efficient [undefz, undefp, undefs]. For robustness, we compute all eigenpairs with $j\in[k]$ in our numerical simulations so we can find the corresponding orthonormal eigenfunctions. As there are local optimizers, the algorithm may converge to nearby optimizers which are close to the initial guess. In this paper, we show the best optimizers only.
@@ 5.1. Minimizers for the disk
Interestingly, $\rho_{k}^{\bigtriangledown}$ is a bang-bang function for all $k\geq 1$.
@@
5.2. Maximizers for the disk
For $k\geq 2$, the maximizer $\rho_{k}^{\triangle}$ in (8) has a non-trivial form shown in Figures 4 and 5. It is the $k$-repetition of an interval that is constant (=1=1) on a first subinterval and is strictly positive and convex on a second subinterval. In the right panels of Figure 4, we identify the Steklov boundary in red. As the maximizers are not bang-bang, the total length of red curves is not necessarily $\pi$ while the constraint $\int_{0}^{2\pi}\rho(\theta)\,d\theta=\pi$ is satisfied.
In Figure 6, we plot $\alpha$ vs. $\alpha\lambda_{2,\alpha}^{\triangle}$. This indicates that, for $k$ even, except in the limit that $\alpha\to 0$, we have that $\lambda_{2,\alpha}^{\triangle}<\frac{2}{\alpha}$, implying that $\lambda_{k,\alpha}^{\triangle}=\frac{k}{2}\lambda_{2,\alpha}^{\triangle}<\frac{k}{\alpha}=\frac{2\pi}{\alpha\left\lvert\Gamma\right\rvert}k$, which shows that the Hersch-Payne-Schiffer inequality [undefo, undefk] is not strict for the extremal weighted Steklov eigenvalue problem (and hence the extremal S-N eigenvalue problem) for $k$ even and $\alpha>0$.
We remark that the oscillations for the maximizing densities, $\rho_{k}^{\triangle}$, are reminiscent of the maximizing domains with oscillatory boundaries for the shape optimization problem studied in [undef].
@@ 5.2. Maximizers for the disk
Figure 4. (Left) Maximizing densities, $\rho_{k}^{\triangle}$, and associated eigenfunctions for (8) with $\Omega=\mathbb{D}$, $\alpha=0.5$, and even $k=2,4,6,8$ (values indicated). Interestingly, the eigenfunctions are constant on intervals where $\rho\neq 1$. (Right) The eigenfunctions are plotted on $\Omega=\mathbb{D}$. The red part of the boundary is where $\rho=1$, i.e., where the Steklov boundary condition is imposed. See Section 5.2.
@@
Figure 5. Solutions for (8) with $\Omega=\mathbb{D}$, $\alpha=0.5$, and $k=1,2,3,4$ (values indicated). In the left hand panels, the maximizing density $\rho_{k}^{\triangle}$ is plotted in blue and the corresponding eigenfunction(s) are plotted in red. For $k$ even, the eigenvalue is simple and the right hand panel shows that the optimality condition (11) is satisfied. For $k$ odd, the eigenvalue has multiplicity 22 and the right hand panel shows that the optimality condition (11) is satisfied for the two eigenfunctions plotted in the left hand panel. See Section 5.2.
@@
(22) $\frac{\lambda_{k,\alpha}^{\bigtriangledown}}{\lambda_{k,\alpha}^{W}}\to\alpha\sigma_{1,\alpha}^{\bigtriangledown}=\alpha\sigma_{1}(\mathcal{I}_{\alpha}^{2})\qquad\qquad\textrm{and}\qquad\qquad\frac{\lambda_{k,\alpha}^{\triangle}}{\lambda_{k,\alpha}^{W}}\to\alpha\lambda_{2,\alpha}^{\triangle}.$
In Figure 6, we plot $\alpha$ vs. $\alpha\lambda_{1,\alpha}^{\bigtriangledown}=\alpha\sigma_{1,\alpha}^{\bigtriangledown}$ and $\alpha\lambda_{2,\alpha}^{\triangle}$. We see that for $\alpha$ small, say $\alpha=0.02$, we have that $\alpha\sigma_{1,\alpha}^{\bigtriangledown}\approx 0.2$ which indicates that $\sigma_{k}(\mathcal{I}_{\alpha}^{k+1})$ is only $20\%$ what Weyl’s law would predict! Furthermore, for $\alpha$ small, we have that $\alpha\lambda_{2,\alpha}^{\triangle}\approx 2$ which indicates that $\lambda_{k,\alpha}^{\triangle}$ is twice what Weyl’s law would predict for large $k$. For other examples of where extremal eigenvalues are a multiplicative factor different than predicted by Weyl’s law, see [undefk, undef, undefr, undefu, undeft].
@@ Appendix A A counterexample to convexity properties of higher eigenvalues
with $0\leq\alpha<\beta$. In [undefn, Theorem 9.1.3], it is claimed that $\rho\mapsto\lambda^{-1}_{k}(\rho)$ is convex for fixed $k\geq 1$, but this is not true for $k\geq 2$ (as the author agreed in personal communication).
@@
(23) $2\tan(\sqrt{\lambda})=-\tan\left(2\sqrt{\lambda}\right)=\frac{-2\tan(\sqrt{\lambda})}{1-\tan^{2}(\sqrt{\lambda})},$
@@
(24) $f_{t}\big|_{t=0,\,\lambda=\eta_{k}}=-\frac{3}{4}\left(-2\tan(2\sqrt{\eta_{k}})+\sqrt{\eta_{k}}\sec^{2}(2\sqrt{\eta_{k}})+\tan(\sqrt{\eta_{k}})-4\sqrt{\eta_{k}}\sec^{2}(\sqrt{\eta_{k}})\right).$
@@
Finally, we remark that convexity/concavity properties for higher eigenvalues of other operators is also misstated, e.g., [undefn, Theorem 8.1.2] for the Schrödinger equation and [undefn, Theorem 10.1.1] for the conductivity equation.